Operations Report – May 8th

Crew 212 Operations Report 08-MAY-2019

SOL: 3

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Oven in the kitchen

Notes on non-nominal systems: The kitchen oven door doesn’t fully close and it was necessary to block it with some tape to use it.

Generator (hours run last night): 12:30

On (last night): 6:30 PM

Off (this morning): 7:00 AM

Solar— SOC 86% at 7:40PM

Diesel Reading – 97%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 270 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 170 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1433682

Toilet tank emptied: Yes

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.1h (the previous value of 92.6h was wrong, the correct one was 91.6h)

Beginning charge: 100%

Ending charge: 80%

Currently charging: Yes

Opportunity rover used: No

Hours: 58.7h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 94.3h

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: An Engineering EVA was done to recover a payload in Spirit’s trunk.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: Today’s EVA was shortened because of the weather.

Summary of internet: Nothing to report

Summary of suits and radios: A major check was run, and each crew member was assigned a functioning radio.

Summary of Hab operations: The oven door doesn’t fully close, so we closed it with tape to cook biscuits. A leak on the lower floor was found after the rain at noon, close to the Engineering Airlock door (right above the propane line).

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: It was depressurized to allow the

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report